This is from a worksheet. I do not understand how to do this at all. I have notes from my instructor but they are not very clear to me about this type of problem. The directions are as follows: "Find all the real number solutions for each equation." Here is my first equation.
x^5-x=0

Answer by scott8148(6628)

Factor, factor, factor...

x^5-x=0 factors into x(x^4-1)=0 ... so x=0 is one solution

x^4-1=0 factors into (difference of 2 squares) (x^2+1)(x^2-1)=0

x^2-1=0 factors into (x+1)(x-1)=0 ... so x=-1 and x=1 are solutions

x^2+1=0 ... x^2=-1 ... no real solutions